AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:14 2025
Author
*
:
Subject
*
:
<p>New run 52 with:</p> <p><strong>Electron cooler setpoint 244.2 V</strong></p> <p>Electron cooler read 243.66 V</p> <p>Collector current 4.3 uA</p> <p>Right arm at 64.60 mm</p> <p>Left arm IN at 36.54 mm</p> <p>Target 2E11 fluctuating</p> <p>Injection ca. 5E7</p> <p> </p> <p><strong>22;00</strong></p> <p>Detectors moved out while HKR plays with the beam injection to improve its stability</p> <p><strong>22;56</strong></p> <p>All system checks good. Detectors still out. Stats in attachment 1</p> <p><strong>23:08</strong></p> <p>HKR still has to call, but I have put detectors back in since they are only operating on injection.</p> <p>Noticed that from time to time the target density still sparks to E12, and this corresponds to when the temperature drops.</p> <p><strong>23:42</strong></p> <p>HKR called back, they found the problem and fixed it. Beam seems more stable, with injections at ca. 5E7</p> <p><strong>23:50</strong></p> <p>DSSD0 high stats, cycled the motors to clear them.</p> <p><strong>00:18</strong></p> <p>All system checks are fine. Attachment 2-3 with stats and a spectrum</p> <p><strong>00:38</strong></p> <p>Spikes in the target got more frequent. Increased temperature to 30.5K and monitor if it helps stabilizing it.</p> <p><strong>01:52</strong></p> <p>All checks OK. Target behaving, density 5-8E11.</p> <p>Attachments 4-5</p> <p><strong>03:10</strong></p> <p>System checks OK.</p> <p>Target consistent</p> <p>Some beam outages which were quickly fixed by HKR</p> <p><strong>05:03</strong></p> <p>System checks OK.</p> <p>Target consistent</p> <p><strong>07:00</strong></p> <p>Lost beam. Detectors moved out</p> <p>Will likely end run once beam is back and change energy. Do not use data past 7 am</p> <p>All system checks OK</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown